OpenStudy (anonymous):

Solve for x: (x – 4)2 = 2

OpenStudy (anonymous):

X = 5

OpenStudy (anonymous):

2x-8=2 x=(2+8):2 x=5

OpenStudy (anonymous):

x = 4 ± sqrt2 x = 4 ± sqrt2 x = 4 ± sqrt2 x = 4 ± sqrt13 those are the answers

OpenStudy (anonymous):

x = 4 ± sqrt2 x = 4 ± 2sqrt13 x = 4 ± 2sqrt2 x = 4 ± sqrt13

OpenStudy (anonymous):

I see You shuould write (x-4)^2

OpenStudy (anonymous):

\[x=4\pm \sqrt{2}\]

OpenStudy (anonymous):

if square you should write ^2 not 2 , we will missunderstand